### v3

@@ 3 Problem Description
where $E\coloneqq\left\{(i,j):i,j\in\mathcal{I}\cup\{d\},\ i\neq j\right\}$. We then define the set of possible travel time delays by
$\Omega\coloneqq\underset{(i,j)\in E}{\bigtimes}B_{i,j}.$
@@
Then, the vector of stochastic travel times $t:\ \Omega\rightarrow\mathbb{N}_{>0}^{|E|}$ is given by
$t(\omega)$ $=t^{\mathrm{det}}+\omega\quad\forall\omega\in\Omega$
@@
for all $(i,j)\in E$. Consistently, we can represent $t(\omega)$ as $t(\omega)=(t_{i,j}(\omega))_{(i,j)\in E}=(t_{i,j}^{\mathrm{det}}+\omega_{i,j})_{(i,j)\in E}$. We assume that events in $\Omega$ are independent, thus $(t_{i,j})_{(i,j)\in E}$ are independent, non-negative random variables with finite support and a known probability distribution $\Xi_{i,j}:\mathbb{N}_{\geq 0}\rightarrow[0,1]$ with $\Xi_{i,j}(t)\coloneqq\mathbb{P}(t^{det}_{i,j}+\omega=t)$ and $\underset{t\in\mathbb{N}_{\geq 0}}{\sum}\Xi_{i,j}(t)=1$. In the following, $t_{i,j}^{\mathrm{det}}$ is also called deterministic or best-case travel time. Additionally, we denote the worst-case scenario of travel times by
$\omega_{\mathrm{max}}=\left(\max\{\omega_{i,j}:\omega_{i,j}\in B_{i,j}\}\right)_{(i,j)\in E}$
@@
If a worker team arrives at task $i$ before its time window opens, it has to wait until $\mathrm{ES}_{i}$ to start the service. Because travel times are stochastic, the start and finish times are random variables, which depend on the worker profile $q$ used to execute the route, the sequence $r$ of tasks, and a realization $\omega\in\Omega$ of travel times. Formally, we define the start and finish times $S_{i}^{r,q}:\Omega\rightarrow\mathbb{N}$ and $F_{i}^{r,q}:\Omega\rightarrow\mathbb{N}$ as
$S_{i}^{r,q}(\omega)$ $\coloneqq\max\{F_{i-1}^{r,q}(\omega)+t_{i-1,i}(\omega),\mathrm{ES}_{i}\},\quad\text{and}$
@@
where $F_{d}^{r,q}(\omega)=\mathrm{tl}^{r,q}$ is equal to the depot leave time of team route $(r,q)$ and i1i-1 is the predecessor of task $i$ along $r$. Furthermore, we denote by $\mathrm{tr}^{r,q}$ the worst-case return time of team route $(r,q)$, which is realized when the worst-case scenario $\omega_{\mathrm{max}}$ occurs.
As the actual travel times are random and hardly predictable beforehand, finishing tasks after the time window closes, i.e., after $\mathrm{LF}_{i}$, is sometimes inevitable. Nevertheless, baggage handling operators often have to guarantee a minimum service level, i.e., the percentage of flights that are not delayed due to improper baggage handling. Moreover, in order to limit the probability of passengers missing their connecting flights, airlines often specify maximum acceptable delays in case of disruptions. Therefore, we allow tasks to be finished after their time window closes, while we limit said violation to at most a fixed amount of time, leading to a new extended latest finish time $\mathrm{LF}_{i}^{\mathrm{e}}\geq\mathrm{LF}_{i}$ for each task $i$. Furthermore, to limit potential financial damages and customer dissatisfaction, we introduce a minimum service level $\alpha\in[0,1]$ and demand that the finish time $F_{i}^{r,q}$ of each task $i$ on each route $r$ must satisfy
@@
Constraint (3), also called chance constraint or service level constraint, limits the probability of delays caused by the baggage handling operator, while constraint (4) guarantees that potential delays do not exceed a prespecified limit. A route $r$ with profile $q$ is called feasible if (3)–(4) are satisfied for all $i\in\mathcal{I}^{r}$. We note that, for planning horizons of up to a few hours, the arrival times of aircraft are typically subject to very little uncertainty as there are no major causes for delay once a plane has departed from its origin airport. Moreover, delays of departing flights, i.e., loading tasks, are either known far in advance or realize shortly before the time window opens. In the former case, delays can be incorporated into the definition of time windows. In the latter case, one is typically unable to readjust the current plan as the loading has already begun and, if said delays were modeled as stochastic time windows, one might obtain overly robust solutions due to the low probability of such events happening. Therefore, the assumption of deterministic time windows is reasonable for short- to medium-term planning horizons.
@@
Each task $i\in\mathcal{I}$ has an assigned weight $w_{i}\geq 0$ that indicates its importance in the flight schedule. This can be, for instance, the number or percentage of passengers that have to reach a connecting flight at the destination airport. The expected cost of a route $r$ with profile $q$ is then given by
@@
where $\mathrm{EF}_{i}\coloneqq\mathrm{ES}_{i}+\min\{p_{i,q}:\ q\in\mathcal{Q}_{i}\}$ is the earliest possible finish time of task $i$ and $P_{i}:(\mathrm{LF}_{i},\mathrm{LF}_{i}^{\mathrm{e}}]\rightarrow\mathbb{R}_{>0}$ is a penalty function. The first part of the objective functions aims to minimize weighted expected finish times in order to have as large safety time buffers as possible to absorb delays in other parts of the baggage handling process. The second part consists of a penalty for delaying flights. Because the probability of passengers missing connecting flights increases superlinearly in the delay of a given baggage handling task, the baggage handling operator aims to keep the variance of delays as low as possible. As the variance can be seen as a quadratic function of the delay, we use a quadratic penalty function
@@
$\mathbb{P}(S^{r,q}_{j}=\tau)=\begin{cases}\begin{aligned} &\underset{z=0}{\overset{\mathrm{ES}_{j}}{\sum}}\mathbb{P}(F^{r,q}_{i}+t_{i,j}=z)&&\text{if}\ \ \tau=\mathrm{ES}_{j}\\ &\mathbb{P}(F^{r,q}_{i}+t_{i,j}=\tau)&&\text{if}\ \ \tau>\mathrm{ES}_{j}\\ &0&&\text{otherwise}\end{aligned}\end{cases}$ (5)
@@
The finish time distribution $F_{j}^{r,q}$ can then be obtained because $F_{j}^{r,q}(\omega)=S_{j}^{r,q}(\omega)+p_{j,q}$ holds by definition for all $\omega\in\Omega$.

### v4

@@ 2 Problem Description
where $E\coloneqq\left\{(i,j):i,j\in\mathcal{I}\cup\{d\},\ i\neq j\right\}$. For each $\mathcal{B}_{k}\in\mathcal{B}$, we then define the set of possible travel time delays by
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Omega^{k}}\coloneqq\underset{(i,j)\in E}{\prod}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}B_{i,j}^{k}}.$
@@
Then, the vector of stochastic travel times ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}t^{k}:\ \Omega^{k}}\rightarrow\mathbb{N}_{>0}^{|E|}$ is given by
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}t^{k}}(\omega)$ $=t^{\mathrm{det}}+\omega\quad\forall\omega\in{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Omega^{k}}$
@@
for all $(i,j)\in E$ and $\mathcal{B}_{k}\in\mathcal{B}$. Consistently, we can represent ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}t^{k}}(\omega)$ as ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}t^{k}}(\omega)=({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}t^{k}}_{i,j}(\omega))_{(i,j)\in E}=(t_{i,j}^{\mathrm{det}}+\omega_{i,j})_{(i,j)\in E}$. Given $k\in\{1,\dots,|\mathcal{B}|$, we assume that events in $\Omega^{k}$ are independent, thus $({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}t^{k}}_{i,j})_{(i,j)\in E}$ are independent, non-negative random variables with finite support and a known probability distribution ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Xi_{i,j}^{k}}:\mathbb{N}_{\geq 0}\rightarrow[0,1]$ with ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Xi^{k}}_{i,j}(t)\coloneqq\mathbb{P}(t^{det}_{i,j}+\omega=t)$ and $\underset{t\in\mathbb{N}_{\geq 0}}{\sum}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Xi^{k}}_{i,j}(t)=1$. In the following, $t_{i,j}^{\mathrm{det}}$ is also called deterministic or best-case travel time. Note that for time bins $k,\bar{k}\in\{1,\dots|\mathcal{B}|\}$ with $k\neq\bar{k}$, $t^{k}_{i,j}$ and $t^{\bar{k}}_{i^{*},j^{*}}$ can be dependent for any $(i,j),(i^{*},j^{*})\in E$. Additionally, given an arbitrary $\gamma\in[0,1]$, we denote the $\gamma$-quantile scenario of distributions $(t_{i,j}^{k})_{(i,j)\in E}$ by
$\omega_{\gamma}^{k}\coloneqq\underset{(i,j)\in E}{\prod}\min\left\{\omega_{i,j}^{k}:\omega_{i,j}^{k}\in\Omega_{i,j}^{k}\ \land\ \mathbb{P}(t_{i,j}^{k}\leq\omega_{i,j}^{k})\geq\gamma\right\}$
Moreover, we define $\omega_{\gamma}\coloneqq\left(\omega_{\gamma}^{k}\right)_{k:\mathcal{B}_{k}\in\mathcal{B}}$ and the worst-case scenario of travel times by
$\omega_{\mathrm{max}}=\left(\max\{\omega_{i,j}^{k}:\omega_{i,j}^{k}\in B^{k}_{i,j}\}\right)_{(i,j)\in E,k=1,\dots,|\mathcal{B}|}$
In the following, the finish times in scenario $\omega_{y}$ are also called $\omega_{\gamma}$-scenario finish times
@@
If a worker team arrives at task $i$ before its time window opens, it has to wait until its earliest start $\mathrm{ES}_{i}$ to start the service. Because travel times are stochastic, the start and finish times are random variables, which depend on the worker profile $q\in\mathcal{Q}$ used to execute the route, the sequence $r$ of tasks, and a realization of travel times. Formally, given a team route $(r,q)\in\mathcal{R}$ and its median finish time $\bar{F}_{i-1}^{r,q}$ at task $i-1\in\mathcal{I}$, we define the start and finish times $S_{i}^{r,q}:\Omega^{\bar{F}_{i-1}^{r,q}}\rightarrow\mathbb{N}$ and $F_{i}^{r,q}:\Omega^{\bar{F}_{i-1}^{r,q}}\rightarrow\mathbb{N}$ at the subsequent task $i\in\mathcal{I}$ as
$S_{i}^{r,q}(\omega)$ $\coloneqq\max\{F_{i-1}^{r,q}(\omega)+t_{i-1,i}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{F}_{i-1}^{r,q}}}(\omega),\mathrm{ES}_{i}\},\quad\text{and}$
@@
where $F_{d}^{r,q}(\omega)=\mathrm{tl}^{r,q}$ is equal to the depot leave time of team route $(r,q)$ and i1i-1 is the predecessor of task $i$ along $r$. Here, $\Omega^{\bar{F}_{i-1}^{r,q}}$ is the distribution of travel times in time bin $\mathcal{B}_{k}\in\mathcal{B}$ with $\bar{F}_{i-1}^{r,q}\in\mathcal{B}_{k}$. Note that the above definition of $S_{i}^{r,q}$ uses the same travel time distribution based on $\bar{F}_{i-1}^{r,q}$ for all events $\omega\in\Omega^{\bar{F}_{i-1}^{r,q}}$ rather than event-dependent distributions based on $F_{i-1}^{r,q}(\omega)$. While this choice poses a slight simplification of reality, it allows for efficient computation of distributions via convolutions and stronger comparability of task sequences. Allowing the current time bin to depend on the realized finish time would require tracking joint distributions over finish times and time bins, leading to a combinatorial explosion in the state space and rendering an exact solution approach as proposed in the subsequent sections intractable. Finally, we note that, in contrast to arbitrary quantiles of finish times, the usage of median finish times in the above context is unbiased towards earlier or later bins and thus reduces the probability of selecting a current time bin that is equal to the true underlying bin only for very few, i.e., outlier events. Furthermore, we denote by $\mathrm{tr}^{r,q}$ return time of team route $(r,q)$, which is realized when case scenario $\omega_{\gamma}$ occurs.
As the actual travel times are random and hardly predictable beforehand, finishing tasks after the time window closes, i.e., after $\mathrm{LF}_{i}$, is sometimes inevitable. Nevertheless, baggage handling operators often have to guarantee a minimum service level, i.e., the percentage of flights that are not delayed due to improper baggage handling. Moreover, in order to limit the probability of passengers missing their connecting flights, airlines often specify maximum acceptable delays in case of disruptions. Therefore, we allow tasks to be finished after their time window closes, while we limit said violation to at most a fixed amount of time, leading to a new extended latest finish time $\mathrm{LF}_{i}^{\mathrm{e}}\geq\mathrm{LF}_{i}$ for each task $i\in\mathcal{I}$. Furthermore, to limit potential financial damages and customer dissatisfaction, we introduce a minimum service level $\alpha\in[0,1]$ and demand that the finish time $F_{i}^{r,q}$ of each task $i\in\mathcal{I}$ on each route $r$ must satisfy
@@
Constraint (3), also called chance constraint or service level constraint, limits the probability of delays caused by the baggage handling operator, while constraint (4) guarantees that potential delays do not exceed a prespecified limit. A route $r$ with profile $q$ is called feasible if (3)–(4) are satisfied for all $i\in\mathcal{I}^{r}$. We note that, for planning horizons of up to a few hours, the arrival times of aircraft are typically subject to very little uncertainty as there are no major causes for delay once a plane has departed from its origin airport. Moreover, delays of departing flights, i.e., loading tasks, are either known far in advance or realize shortly before the time window opens. In the former case, delays can be incorporated into the definition of time windows. In the latter case, one is typically unable to readjust the current plan as the loading has already begun and, if said delays were modeled as stochastic time windows, one might obtain overly robust solutions due to the low probability of such events happening. Therefore, the assumption of deterministic time windows is reasonable for short- to medium-term planning horizons. Finally, earliest start times for incoming aircraft and latest finish time for outgoing aircraft are determined by the arrival time at the gate and the scheduled departure time, respectively. Latest finish times for incoming aircraft are defined based on baggage transfer time requirements alongside maximum baggage claim durations, while earliest start times for outgoing flights depend on the aircraft’s earliest readiness time.
@@
Each task $i\in\mathcal{I}$ has an assigned weight $w_{i}\geq 0$ that indicates its importance in the flight schedule. This can be, for instance, the number or percentage of passengers that have to reach a connecting flight at the destination airport. The expected cost of a team route $(r,q)\in\mathcal{R}$) is then given by
@@
where $\mathrm{EF}_{i}\coloneqq\mathrm{ES}_{i}+\min\{p_{i,q}:\ q\in\mathcal{Q}_{i}\}$ is the earliest possible finish time of task $i\in\mathcal{I}$ and $P_{i}:(\mathrm{LF}_{i},\mathrm{LF}_{i}^{\mathrm{e}}]\rightarrow\mathbb{R}_{>0}$ is a penalty function. The first part of the objective functions aims to minimize weighted expected finish times in order to have as large safety time buffers as possible to absorb delays in other parts of the baggage handling process. The second part consists of a penalty for delaying flights. Because the probability of passengers missing connecting flights increases superlinearly in the delay of a given baggage handling task, the baggage handling operator aims to keep the variance of delays as low as possible. As the variance can be seen as a quadratic function of the delay, we use a quadratic penalty function
@@
$\mathbb{P}(S^{r,q}_{j}=\tau)=\begin{cases}\begin{aligned} &\underset{z=0}{\overset{\mathrm{ES}_{j}}{\sum}}\mathbb{P}(F^{r,q}_{i}+t_{i,j}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{F}_{i}^{r,q}}}=z)&&\text{if}\ \ \tau=\mathrm{ES}_{j}\\ &\mathbb{P}(F^{r,q}_{i}+t_{i,j}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{F}_{i}^{r,q}}}=\tau)&&\text{if}\ \ \tau>\mathrm{ES}_{j}\\ &0&&\text{otherwise}\end{aligned}\end{cases}$ (5)
@@
The finish time distribution $F_{j}^{r,q}$ can then be obtained because $F_{j}^{r,q}(\omega)=S_{j}^{r,q}(\omega)+p_{j,q}$ holds by definition for all $\omega\in\Omega^{k}$ and $k=1,\dots,|B|$.
3 Literature Review
The problem considered in this paper can be seen as a variant of the technician scheduling and routing problem (TSRP). TSRPs typically consist of a routing part that can be seen as a vehicle routing problem (VRP) and a scheduling or team formation part. The solution approach proposed in the following is an exact approach and works with arbitrary distributional structures. Section 3.1 focuses on literature dealing with stochastic VRPs in a general context, with a particular emphasis on assumptions on the distributional structure of random variables and the assumption of independence. Section 3.2 considers publications dealing with stochastic TSRP variants.
3.1 Literature on Stochastic Vehicle Routing
Recently, stochastic formulations of the vehicle routing problem with time windows (VRPTW) have seen a noticeable interest increase. In the following, we provide an overview of literature dealing with stochastic variants of the VRPTW. Because distributional assumptions and assumptions regarding independence of random variables have a significant impact on the computational tractability of exact approaches, as well as the practical applicability of results, we especially focus on these two aspects. oyola2018stochastic summarize the most relevant literature on common types of stochastic VRPs. A taxonomy and overview of the space of VRPs can be found in eksioglu2009vehicle.
Generally, random variables in the stochastic VRPTW are assumed to be independent as stochastic dependency drastically complicates the calculation of distributions and their moments. While some publications focus on the VRPTW with stochastic demands (see lee2012robust, zhang2016time), stochastic travel and service times are far more common. Stochasticity can be addressed in different ways. A common approach is to include chance constraints, which limit the probability of violating time windows. While errico2018vehicle connect chance constraints to the simultaneous satisfaction of all time windows, ehmke2015ensuring as well as li2010vehicle limit the probability of violating the time window of each customer individually.
In vehicle routing problems, the distributions of start times at a given customer depend on the sequence of customers previously visited by the same vehicle and the distributions of travel and service times along the route. Furthermore, the computational effort required to calculate said distributions heavily depends on the type of time windows. tacs2014vehicle assume that time windows are soft and travel times are independent, gamma distributed random variables. The authors propose an exact calculation of arrival time distributions by convolving finish time and travel time distributions and solve the problem using a Branch-Price-and-Cut approach. Because the convolution of gamma distributed random variables is gamma distributed, finish time distributions are again gamma distributed with known shape and scale parameter, which significantly simplifies their calculation.
If time windows are assumed to be hard, arrival time distributions need to be truncated at the start of each time window. Hence, distributional properties, such as moments or the existence of a closed-form expression, are not propagated along vehicle routes. Therefore, distributions can only be obtained by explicitly calculating the convolution of two distributions of arbitrary shapes, which is often not computationally tractable in reasonable time. errico2018vehicle assume service times to be stochastic, discretely distributed random variables and calculate start time distributions exactly. They use a Branch-Price-and-Cut algorithm to solve the problem and report computational results for four different variants of triangular distributions. miranda2016vehicle assume normally distributed service and travel times. They approximate service and travel time distributions by discretizing them on a finite set of values, whose count and values depend on the underlying distribution’s moments. This allows for an approximation of the finish time distribution of a given task by convolving the finish time distribution of the previous task with the aforementioned approximations of travel and service time distributions. They further improve their approach in miranda_revised using improved lower bounds for the approximation intervals of service and travel times. Nevertheless, distributions are approximated, and thus the resulting algorithm is not exact. ehmke2015ensuring show that the first and second moment of start time distributions can be calculated with little effort if both travel times and start times at the previous customer are normally distributed. While the former is usually not the case, the authors present computational results that indicate that their approach provides suitable approximations for normal, shifted gamma, and shifted exponential travel time distributions. li2010vehicle use stochastic simulation to derive estimates for start time distributions and solve the problem using tabu search. To the best of our knowledge, there is no publication that proposes an exact solution approach for medium- to large-sized VRPTW instances with stochastic travel times of arbitrary distributional shape.
We note that the distribution of travel times on airport aprons often does not follow standard distributional structures such as normal or gamma distributions. Furthermore, distributions evolve over time and thus change both their moments and their shape. Because this impedes the exploitation of distributional structures as done in the aforecited papers, the aim of this paper is to develop an exact solution approach that is able to handle travel time distributions of arbitrary shape, while also allowing for the consideration of temporal dependency of distributions. To the best of our knowledge, the problem at hand has not been considered under the assumption of time-dependent travel times of arbitrary distributional shapes before. From the point of view of distributional assumptions and usage of chance constraints
, our considerations can be seen as a combination and extension of errico2018vehicle and li2010vehicle, as we do not assume the knowledge of any distributional structures and calculate start time distributions exactly but interpret chance constraints as a non-route-interdependent property.
3.2 Literature on the stochastic TSRP
Due to the lack of standardized model formulations, several publications (cp. pereira2020multiperiod, cakirgil2020integrated, li2005manpower) consider problems that are similar to the problem at hand. However, all of the aforementioned publications either make restrictive assumptions regarding team formation possibilities, such as the impossibility of downgrading or the existence of a single mode, or they focus on heuristic solution methods to solve medium- and large-sized instances. For an overview of further literature dealing with deterministic variants of the TSRP relevant to this study, we refer to dallolio_temp.
While there is plenty of literature regarding the deterministic TSRP, very little research has been done on stochastic problem variants. souyris2013robust propose a robust formulation for the TSRP. Workers are assumed to be homogeneous and service times are stochastic. Each task must be started before a fixed deadline with a given probability. Moreover, each task must be executed by a single technician, thus team formation is not part of the model. The authors propose three different bounded uncertainty sets that limit the total service time delays per technician or client, respectively. The objective is to minimize the worst-case total delay and travel time. A branch-and-price approach is proposed and applied to real-world instances with 41 customers and 15 technicians.
yuan2015branch address the TSRP in the context of health care workers and home health care services. Each task must be executed by a single worker. Furthermore, it is possible to downgrade workers to lower skill levels. Service times are assumed to be stochastic with a known probability distribution. The goal is to minimize the expected total travel costs, fixed costs of caregivers, service costs, and penalties for late arrival at customers. The authors use a branch-and-price approach to solve the problem exactly. Computational experiments assuming uniformly distributed service times and 25 to 50 customers indicate that the approach is able to provide very good results for small-sized instances. While the objective function and interpretation of stochasticity considered by yuan2015branch is different from ours and several properties, such as the possibility of executing tasks using different modes and skill compositions, are missing, their solution approach exhibits similarities with the one proposed in the following and can be applied to the problem at hand. In Section 6.2, we compare both methodologies on a set of test instances.
binart20162 consider a TSRP with mandatory and optional tasks, where time windows are hard and fixed, and predefined teams must be used to execute tasks. Although not explicitly done, their modeling framework would allow considering multiple skill levels, downgrading, and multiple modes for single tasks. They assume that travel and service times are discretely distributed according to triangular distributions. Furthermore, the time windows of each task must be satisfied with a given probability. A 2-stage approach is proposed to heuristically solve the problem. First, a feasible skeleton solution is obtained using a generic MIP solver by optimally covering all mandatory tasks. Second, the first-stage solution is refined by inserting optional customers into the existing routes such that time window restrictions are not violated. Computational results on instances with 5 to 9 mandatory and 30 to 50 optional tasks indicate that the approach can yield good results in most cases. The problem considered by binart20162 can be seen as a special case of the problem at hand where all technicians have the same skill level, each team formation represents one technician, and each task can be executed by each team formation. While the objective function and interpretation of stochasticity differ from ours, the models and solution approach proposed in the following can be applied to their problem.
Our work shows several similarities with binart20162 and yuan2015branch, such as the stochasticity of travel times, the usage of exact solution methods, and the incorporation of several team properties such as multiple skill levels and modes. At the same time, our interpretation of service levels at individual tasks rather than entire solutions, the lack of assumptions on distributional properties, and the exact calculation of arrival times separate our considerations from previous approaches.
The following considerations are widely based on dallolio_temp, which consider the problem of forming and routing worker teams for baggage handling tasks in a deterministic setting. We extend their deterministic problem description by assuming stochastic travel times. Furthermore, we advance their proposed solution algorithm by introducing cutting planes and additional branching strategies, improving the runtime efficiency of the pricing step, and describing an efficient way to handle infeasible solutions that is based on dynamically switching between two master problem formulations.